# Manga109Dialog: A Large-scale Dialogue Dataset for Comics Speaker Detection

### v1

@@ Abstract
The expanding market for e-comics has spurred interest in the development of automated methods to analyze comics. For further understanding of comics, an automated approach is needed to link text in comics to characters speaking the words. Comics speaker detection research has practical applications, such as automatic character assignment for audiobooks, automatic translation according to characters’ personalities, and inference of character relationships and stories. To deal with the problem of insufficient speaker-to-text annotations, we created a new annotation dataset Manga109Dialog [^1] Dataset and code are available at https://github.com/manga109/public-annotations. based on Manga109. Manga109Dialog is the world’s largest comics speaker annotation dataset, containing 132,692 speaker-to-text pairs. We further divided our dataset into different levels by prediction difficulties to evaluate speaker detection methods more appropriately. Unlike existing methods mainly based on distances, we propose a deep learning-based method using scene graph generation models. Due to the unique features of comics, we enhance the performance of our proposed model by considering the frame reading order. We conducted experiments using Manga109Dialog and other datasets. Experimental results demonstrate that our scene-graph-based approach outperforms existing methods, achieving a prediction accuracy of over 75%.
@@ 1 Introduction
The market for e-comics has been expanding rapidly with the popularization of digital devices. In 2022, e-comics made up 66.2% of the Japanese comics market shuppankagaku2023 . The development of e-comics has led to increased interest in automated methods to analyses comics. To support these techniques, reliable large-scale dialogue datasets are required for improved computational understanding of comics. Additionally, automated methods to detect characters to whom text is attributed are necessary for effective speaker detection.
Figure 1: An example of speaker detection. ©Akamatsu Ken.
@@
An example of speaker detection is shown in Figure 1. Given an image from comics, the system can detect characters, text, and frame regions, and can then automatically predict the character to which the text is attributed. This research can be applied to various tasks, such as automatic character assignment for text-to-speech reading, automatic translation according to characters’ personalities, inference of character relationships and stories, and automated generation of scenarios.
Due to the lack of dialogue datasets and some limitations of existing annotations, we constructed more accurate and rich annotations, Manga109Dialog. To the best of our knowledge, this is the largest comics dialogue dataset ever created. We divided the test set into two subsets based on the prediction difficulty.
@@
To evaluate the reliability and accuracy of Manga109Dialog, we conducted speaker detection experiments on it. Previous approaches were primarily based on the rule that the character nearest to a given text is most likely to be the speaker rigaud2015speech ; abe2020 . However, due to the unique and complex structure of comics, they may have difficulty making correct predictions in some complex cases. As an example, consider the text at bottom right in Figure 1. If we chose the character nearest to the text as the speaker, the boy would be incorrectly predicted as the speaker rather than the girl. Therefore, the relationship between characters and text should be considered to make a correct prediction.
@@
To address this challenge, we propose an innovative deep learning-based approach. We leverage scene graphs, which are widely recognized as one of the most popular methods for describing visual relationships. Scene graph generation (SGG) is the task of detecting objects and their relationships in an image. Since SGG models have performed well on real-world datasets, such as the Visual Genome dataset krishna2017visual , we propose utilizing SGG models for the task of comics speaker detection. Along with the general framework of SGG models, we introduce frame information because we find that the speaker tends to appear in the same frame as the text.
@@
Traditional metrics used to evaluate SGG lack suitability because the amount of speaker-to-text pairs varies from page to page. Besides, prediction results using traditional metrics cannot be compared to those of rule-based methods. Therefore, we present an evaluation metric well suited for this specific task, and demonstrate its superiority over conventional approaches in experimental results on Manga109Dialog. The contributions are summarized as follows.
@@
  • We constructed an annotation dataset of associations between speakers and texts. This is the largest comics dialogue dataset in the world, containing over 130,000 speaker-to-text pairs.
@@
  • We propose a deep learning-based approach for comics speaker detection using SGG models. We enhance the results by introducing frame information in the relationship prediction stage.
  • We introduce a new evaluation metric and a new standard for data division, that is, dividing the dataset according to prediction difficulty.
@@
Experimental results showed that Manga109Dialog provides a challenging yet realistic benchmark for comics speaker detection, and that our proposed approach exhibited 5% better performance compared to conventional rule-based methods. This research aimed to show to what extent speaker detection can be achieved solely with visual information, providing a benchmark for incorporating NLP in future research.
@@ 2.1 Comics Speaker Dataset
Among the many comics available in different forms, only a few can be used for academic research due to copyright issues. eBDtheque guerin2013ebdtheque is a dataset of comics available for direct use. It comprises 100 images from French, American, and Japanese comics, containing 1,550 characters, 1,092 speech balloons, and 887 speaker-to-text pairs. Moreover, a large-scale manga dataset called Manga109 aizawa2020building contains 109 Japanese comics, including 21,142 images, 2,979 characters, and 147,918 texts. Given the lack of available comics datasets associated with speakers and comics texts, Abe et al. abe2020 constructed a dataset based on Manga109 with a total of 121,364 annotated speaker-to-text pairs.
@@ 2.2 Comics Speaker Detection
Rule-Based Methods:
Rigaud et al. rigaud2015speech proposed early rule-based speaker prediction method using distance calculation between speech balloon centroids and character centroids. The character with the shortest distance was assumed to be the speaker. They performed experiments on eBDtheque and obtained an accuracy of 78.58%. In the study by Abe et al. on Manga109, four types of information were used abe2020 , including the text-character distance, the characters in the same frame with the text, the direction of the tail of the speech balloon, and language style (such as first-person utterances). Their approach experimentally obtained an accuracy of 70.7%.
Deep Learning-Based Methods:
In recent years, deep learning-based methods have also been developed. Yamamoto et al. yamamoto2018 used features extracted from the input images and metadata to calculate scores for each text-character pair, while other studies explored time-series learning nakagawa2019 and Natural Language Processing yamaguchi2021 . The results of these studies on Manga109 all demonstrate the potential of learning-based methods in speaker prediction. However, their results were only based on a part of the dataset and lacked implementation details. Therefore, a large dataset is needed to evaluate these models under a unified standard.
@@ 2.3 Scene Graph Generation (SGG)
Scene graphs were introduced by Johnson et al. johnson2015image as a data structure for describing objects in a scene. To generate a scene graph, an SGG model first detects object regions and their categories, then identifies relationships between objects. Since scene graphs provide rich semantic features, they can distinguish images or videos more accurately and describe them more precisely. Scene graphs have demonstrated success in various visual tasks, including image retrieval, image captioning, and visual question answering johnson2015image ; yao2018exploring ; tang2019learning .
@@ 3 Manga109Dialog
Abe et al. abe2020 first created a comics speaker dataset based on the entire Manga109 dataset. Although their study was pioneering and impressive, this dataset still involves some notable limitations. First, they linked the name of the character to the text instead of the bounding box of the speaker. Therefore, we cannot specify a specific bounding box if the same character appears multiple times on a single page. Besides, since annotators prefer making annotations of simple relationships, some more complicated cases are not included in this dataset.
Aiming to produce a more complete and applicable dataset for speaker detection in comics, we constructed a new dataset, Manga109Dialog. Instead of connecting the character name to the text, we connect the bounding box of the speaker to the text. The visual differences between Abe’s dataset and Manga109Dialog are shown as Figure 2.
@@ 3.1 Dataset Construction
In this section, we show the details of how we constructed our dataset. Our annotations were mainly based on the following rules.
@@
  • If more than one character is on the page, the bounding box of the speaker in the same frame as the text is linked in priority. If the speaker is not in the same frame, it is determined by the reading order (from top to bottom, right to left). The texts in the last frame of the page, however, are linked to the character in the second-to-last frame.
@@
Although we make annotations on Manga109, not all texts are our annotation targets. For example, titles and descriptions were not annotated. Texts designed as “Narration” or with unknown speakers were also not annotated. Furthermore, in contrast to Abe’s dataset, when the speaker was not on the page, such as texts from phone conversations or letters, we did not annotate them even if we knew the speaker. More details and examples of annotation rules are shown in the supplementary material.
(a) Abe’s annotations (character-to-text).
(b) Manga109Dialog (speaker-to-text).
Figure 2: Our dataset allows us to accurately determine whether the text on the right side is spoken by the character in the first frame or the character in the second frame, which Abe’s annotations were unable to do. ©Arai Satoshi.
Table 1: Statistics for two datasets.
@@
Dataset Annotated images Texts Easy Hard Total Pairs / page
Manga109Dialog 9,904 147,887 111,959 20,733 132,692 6.70
Abe’s annotations 9,830 147,887 - - 121,291 6.17
@@ 3.2 Data Analysis
Figure 3: An example of annotations on three prediction difficulties. ©Kurita Riku.
We outsourced the annotation of our dataset to a company with expert annotators. After they completed the task, the first author confirmed and corrected their annotations. Creating the annotations took approximately three months. An overview of Manga109Dialog and Abe’s annotation dataset is shown in Table 1. In order to compare the two datasets at the same level, we did an additional step to determine the bounding boxes of speakers. For cases where the same character appears multiple times on a page, we preferred the box in the same frame as the text; otherwise, we would choose the nearest one. From Table 1, we can find Abe annotated 9,830 images of Manga109, containing an average of 6.17 speaker-to-text pairs per page, whereas we annotated all pages. In Manga109Dialog, 9,904 images include speaker-to-text pairs, with 6.70 pairs per page.
Moreover, we divided our annotations into Easy and Hard by the difficulty of prediction. If the speaker is in the same frame as the text, the text is considered Easy; otherwise, it is considered Hard. Total contains all pairs of speakers and texts. An example of annotations of varying difficulties is given in Figure 3. We performed experiments under these three difficulties to better evaluate our dataset in Section 5.3.
4 Approach
Here, we present our proposed scene-graph-based approach for comics speaker prediction using Manga109Dialog.
4.1 Problem setting
Let us first define our problem setting. Given an input image $I$, our task is twofold. First, we localize character regions and text regions. Next, for all combinations of characters and texts, we calculate the relationship score of each combination. The character with a higher score is more likely to be the speaker of the text. An example is shown in Figure 4. We represent each region as a bounding box $\mathbf{b}=[x,y,w,h]^{\top}\in\mathbb{R}^{4}$. For each region, we predict an object label $l\in\{\mathtt{character},\mathtt{text},\mathtt{background}\}$. For the combination of character $i$ and text $j$, the relationship score between $\mathbf{b}_{i}$ and $\mathbf{b}_{j}$ is represented as $r_{i\to j}\in\mathbb{R}$. This problem setting is equivalent to a standard SGG task, where the number of object classes is two (character or text) and the number of relationship labels is one (speak). In addition, we provide a frame detector and a frame order estimator. We can predict a bounding box of a frame; $\mathbf{p}=[x,y,w,h]^{\top}\in\mathbb{R}^{4}$. We can also obtain frame reading orders. We later use this information to enhance our system.
@@ 4.2 Speaker Detection via SGG
Introducing our approach in detail, we follow the pipeline of MOTIFS zellers2018neural , the most representative SGG model. Its process can be divided into three stages. The first stage is object detection. Given an image $I$, an object detector $f_{\mathrm{detect}}$ outputs a set of $N$ tuples.
@@
Each tuple consists of a bounding box $\mathbf{b}_{i}\in\mathbb{R}^{4}$, a feature vector $\mathbf{f}_{i}\in\mathbb{R}^{4096}$, and a vector of class label probabilities $\mathbf{l}_{i}\in\mathbb{R}^{3}$. The feature vector $\mathbf{f}_{i}$ encodes the visual information of the region $\mathbf{b}_{i}$. $\mathbf{l}_{i}$ represents the probability that the label is $\{\mathtt{character},\mathtt{text},\mathtt{background}\}$ respectively. Unlike MOTIFS, we use a Faster-RCNN ren2015faster with a ResNeXt-101-FPN lin2017feature ; xie2017aggregated backbone for this step.
@@
Here, $\mathbf{d}_{i}\in\mathbb{R}^{512}$ is an enhanced representation for $\mathbf{b}_{i}$. Following MOTIFS, we adopt bidirectional LSTM models hochreiter1997long for this step. An additional LSTM is used to predict the object label $l_{i}$ from $\mathbf{l}_{i}$.
In the last stage, we predict the relationship score for each combination of all $N$ bounding boxes. Therefore, there are $N^{2}$ types of predictions. For example, the score of $\mathbf{b}_{i}$ and $\mathbf{b}_{j}$ ($r_{i\to j}\in\mathbb{R}$) is computed as
@@
The function $g$ inputs the RoI feature of the union box $\textbf{f}_{i,j}$ and enhanced features $\mathbf{d}_{i},\mathbf{d}_{j}$ obtained from the previous stage and outputs a score. Here, $g$ consists of simple learnable matrices and a Softmax predicate classifier. Note that the feature vector $\textbf{f}_{i,j}$ is extracted from bounding boxes $\mathbf{b}_{i},\mathbf{b}_{j}$ and their feature vectors $\mathbf{f}_{i},\mathbf{f}_{j}$. The weight function $w(i,j)\in\mathbb{R}$ can be any function. If $w(i,j)=1$, the entire pipeline is the same as that of MOTIFS.
Figure 4: Framework of the proposed method. ©Arai Satoshi.
@@ 4.3 Use of Reading Order
Due to the unique feature of comics that the speaker is more likely to appear in the same or next frame that the text belongs to, we add frame information to the SGG model. To do so, we train another Faster-RCNN model to detect the frame regions from the input image. Following Kovanen and Ikuta’s study kovanen2015layered ; ikuta2023 , we can then estimate the reading order of the frames. Given a set of frame bounding boxes, we first decide whether they are horizontally divisible into two parts. If so, we split them and repeat the first step for each set until we cannot horizontally split them. Then, we decide whether each set is vertically divisible into two parts. If so, we do so and repeat the first step. When the set of frames cannot be divided horizontally and vertically, we number the frames according to reading order (from top to bottom, from right to left). As shown in Figure 4, we can obtain the result as the form of $\mathbf{p}_{1},\mathbf{p}_{2},\dots$, where $\mathbf{p}_{i}\in\mathbb{R}^{4}$ is a bounding box of the $i$-th frame. We also show an another example of frame detection and reading order estimation results in the supplementary material.
@@
Here, $k_{i}$ represents the reading order of the frame $\mathbf{p}$ to which $\mathbf{b}_{i}$ belongs. The closer the reading orders of $\mathbf{b}_{i}$, $\mathbf{b}_{j}$ are, the higher the score of $w(i,j)$. Therefore, we can make effect the prediction output by introducing frame orders and changing relationship scores.
To summarize, the end-to-end pipeline we designed is as follows. Given $I$, we obtain $N$ tuples $(\mathbf{b}_{i},\mathbf{f}_{i},\mathbf{l}_{i},\mathbf{d}_{i})$ by running Eq. 1 and 2. In addition, we obtain frame information $k_{i}$ through our frame detector. We then compute $r_{i\to j}$ for all $N^{2}$ possible combinations by Eq. 3. From the $N^{2}$ scores, we select the top $K$ items. Here, $K$ is a hyper-parameter. We discuss how to select $K$ in Section 4.4. After this, the model can generate a scene graph, consisting of object labels, their bounding boxes, and speaker-to-text pairs.

### v2

@@ Abstract
The expanding market for e-comics has driven the development of automated methods for analyzing comics. To enhance the machine’s understanding of comics, an automated method is essential for linking text in comics to characters that speak those words. In this study, we developed Manga109Dialog [^1] Dataset and code are available at https://github.com/liyingxuan1012/Manga109Dialog., which is the world’s largest speaker-to-text annotation dataset for comics, containing 132,692 pairs. We proposed a novel deep learning-based method using scene graph generation models. To tailor the unique features of comics, we enhanced the performance by considering the frame reading order. Our experiments with Manga109Dialog show that our scene-graph-based approach outperforms existing methods, achieving a prediction accuracy of over 75%, thus establishing a robust benchmark for speaker detection in comics.
Index Terms—  Dataset, Comic, Scene graph generation
@@ 1 Introduction
The market for e-comics has been expanding rapidly with the popularization of digital devices. In 2022, e-comics made up 66.2% of the Japanese comics market [1]. The development of e-comics has led to increased interest in automated methods to analyses comics. To support these techniques, reliable large-scale dialogue datasets are required for improved computational understanding of comics. Additionally, automated methods to detect characters to whom text is attributed are necessary for effective speaker detection.
Fig. 1: An example of speaker detection. Courtesy of Akamatsu Ken.
@@
An example of speaker detection is shown in Fig. 1. Given an image from comics, the system can detect characters, text, and frame regions, and can then automatically predict the character to which the text is attributed. This research can be applied to various tasks, such as automatic character assignment for text-to-speech reading, automatic translation according to characters’ personalities, inference of character relationships and stories.
@@
To address the significant gap in reliable speaker-to-text annotation datasets for comics, we developed Manga109Dialog. To the best of our knowledge, this is the largest comics dialogue dataset ever created. In order to validate the reliability and accuracy of Manga109Dialog, we conducted comprehensive speaker detection experiments on it. Since previous approaches were primarily based on the rule that the character nearest to a given text is most likely to be the speaker [2, 3], they may have difficulty making correct predictions in some complex cases. As an example, consider the text at bottom right in Fig. 1. If we chose the character nearest to the text as the speaker, the boy would be incorrectly predicted as the speaker rather than the girl. Therefore, the relationship between characters and text should be considered to make a correct prediction.
@@
To address this challenge, we propose a novel deep learning-based approach. We leverage scene graphs, a powerful tool for describing visual relationships. The task of scene graph generation (SGG) involves detecting objects and their relationships within an image. Given the successful application of SGG models in analyzing real-world datasets, such as the Visual Genome dataset [4], we propose their adaptation for comics speaker detection. In addition to employing the standard framework of SGG models, we incorporate frame information, as we find that the speaker tends to appear in the same frame as the text.
@@
Traditional metrics used to evaluate SGG lack suitability because the amount of speaker-to-text pairs varies from page to page. Therefore, we present an evaluation metric well suited for this specific task.
The contributions are summarized as follows.
@@
  • We constructed Manga109Dialog, an annotation dataset of associations between speakers and texts. This is the largest comics dialogue dataset in the world.
@@
  • We proposed a deep learning-based approach for comics speaker detection using SGG models. We enhance the results by introducing frame information in the relationship prediction stage.
  • We established a new benchmark for evaluation, setting a standard for future research in this domain.
@@
Experimental results demonstrated that Manga109Dialog provides a challenging yet realistic benchmark for comics speaker detection. Moreover, our proposed approach outperformed conventional rule-based methods by 5%, highlighting its effectiveness in complex cases.
@@ 2.1 Comics Speaker Dataset
Given the scarcity of available comics speaker datasets, Abe et al. [3] first constructed a dataset based on Manga109 [5], which contains 109 Japanese comics and annotations for both character and text regions. Although Abe’s work is pioneering, the dataset has some notable limitations. First, it does not have explicit links between the text and the character in the comics – it only adds the character name to the text without specifying the bounding box of the character, which poses a challenge when the same character appears multiple times on a page. Additionally, the tendency to annotate simpler relationships results in exclusion of more complex cases. To address these shortcomings, we developed Manga109Dialog, employing a unique annotation standard to provide a more comprehensive and applicable dataset.
@@ 2.2 Comics Speaker Detection
Traditional approaches in comics speaker prediction typically focus on predicting the relationships between characters and text without detecting specific object regions. These methods, primarily rule-based [2, 3], determine the speaker by calculating the distances between text and characters, which often makes them struggle in complex cases. Recently, there has been a shift towards deep learning-based approaches, such as extracting features from objects or employing time-series learning techniques [6, 7]. Although they demonstrated the potential of deep learning-based approaches, they were limited by using only partial datasets and showed suboptimal performance.
This gap underscores the need for a complete dataset and a unified standard for evaluation. In our study, we introduce object detection into the speaker prediction evaluation, aligning more closely with real-world applications. We also develop a comprehensive benchmark for this task.
@@ 2.3 Scene Graph Generation (SGG)
Scene graph is a data structure for describing objects in a scene. Since scene graphs provide rich semantic features, they can distinguish images or videos more accurately and describe them more precisely. Scene graphs have demonstrated success in various visual tasks, including image retrieval, image captioning, and visual question answering [8, 9, 10].
@@ 3 Manga109Dialog
Fig. 2: Annotation rules. Courtesy of Inokuma Shinobu.
@@ 3.1 Dataset Construction
In this section, we show the details of how we constructed Manga109Dialog. Aiming to produce a more complete and applicable dataset for speaker detection in comics, we created our dataset in a distinct annotation standard. Instead of connecting the character name to the text, we connect the bounding box of the speaker to the text. The annotations were mainly based on the following rules.
@@
  • If more than one character is on the page, the bounding box of the speaker in the same frame as the text is linked in priority. If the speaker is not in the same frame, it is determined by the reading order (from top to bottom, right to left), as shown in Fig. 2.
@@
More details and examples of annotation rules are shown in the supplementary material.
Table 1: An overview of Manga109Dialog.
Item Number of elements
Images 10,602
Annotated Images 9,904
Texts 147,887
@@
      Easy 111,959
      Hard 20,733
      Total 132,692
Pairs / Page (Total) 6.70
@@ 3.2 Data Analysis
Creating the annotations for Manga109Dialog took approximately three months. An overview of the dataset is presented in Table 1. Manga109Dialog comprises 9,904 images featuring speaker-to-text pairs, averaging 6.70 pairs per page.
We further divided our annotations into Easy and Hard by prediction difficulties. If the speaker is in the same frame as the text, the text is considered Easy; otherwise, it is considered Hard. Total contains all pairs of speakers and texts. An example of annotations of varying difficulties is given in Fig. 3. We conducted experiments under these three difficulties to evaluate speaker detection methods more appropriately.
4 Approach
4.1 Problem Setting
Before introducing our proposed scene-graph-based approach, let us first define our problem setting. Unlike existing studies on comics speaker prediction, we incorporate the steps for detecting and identifying object regions.
Given an input image $I$, our task is twofold. First, we localize character regions and text regions. Then, for all combinations of characters and texts, we calculate the relationship score of each combination. The character with a higher score is more likely to be the speaker.
An example of the pipeline is shown in Fig. 4. We represent each region as a bounding box $\mathbf{b}=[x,y,w,h]^{\top}\in\mathbb{R}^{4}$. For each region, we predict an object label $l\in\{\mathtt{character},\mathtt{text},\mathtt{background}\}$. For the combination of character $i$ and text $j$, the relationship score between $\mathbf{b}_{i}$ and $\mathbf{b}_{j}$ is represented as $r_{i\to j}\in\mathbb{R}$. This problem setting is equivalent to a standard SGG task, where the number of object classes is two (character or text) and the number of relationship labels is one (speak). In addition, we develop a frame detector and a frame order estimator to detect frame bounding boxes ($\mathbf{p}=[x,y,w,h]^{\top}\in\mathbb{R}^{4}$) and determine their reading order. We later use this information to enhance our system.
@@ 4.1 Problem Setting
Fig. 3: An example of annotations on three prediction difficulties. Courtesy of Kurita Riku.
Fig. 4: Framework of the proposed method. Courtesy of Arai Satoshi.
@@ 4.2 Speaker Detection via SGG
Our approach is based on the MOTIFS pipeline [11, 12], the most representative SGG model. Its process is divided into three stages. The first stage is object detection. Given an image $I$, an object detector $f_{\mathrm{detect}}$ outputs a set of $N$ tuples.
@@
Each tuple consists of a bounding box $\mathbf{b}_{i}\in\mathbb{R}^{4}$, a feature vector $\mathbf{f}_{i}\in\mathbb{R}^{4096}$, and a vector of class label probabilities $\mathbf{l}_{i}\in\mathbb{R}^{3}$. Unlike MOTIFS, we employ a Faster-RCNN [13] with a ResNeXt-101-FPN [14, 15] backbone for this step.
@@
Here, $\mathbf{d}_{i}\in\mathbb{R}^{512}$ is an enhanced representation for $\mathbf{b}_{i}$. We adopt bidirectional LSTM models [16] for this step. An additional LSTM is used to predict the object label $l_{i}$ from $\mathbf{l}_{i}$.
In the last stage, we calculate the relationship score for each pair among the $N$ bounding boxes, resulting in $N^{2}$ possible combinations. For example, the score of $\mathbf{b}_{i}$ and $\mathbf{b}_{j}$ ($r_{i\to j}\in\mathbb{R}$) is computed as
@@
The function $g$ inputs the union box’s RoI feature $\textbf{f}_{i,j}$ and enhanced features $\mathbf{d}_{i},\mathbf{d}_{j}$ and outputs a score. Here, $g$ consists of simple learnable matrices and a Softmax predicate classifier. The feature vector $\textbf{f}_{i,j}$ is extracted from bounding boxes $\mathbf{b}_{i},\mathbf{b}_{j}$ and their feature vectors $\mathbf{f}_{i},\mathbf{f}_{j}$. The weight function $w(i,j)\in\mathbb{R}$ can be any function. If $w(i,j)=1$, the entire pipeline is the same as that of MOTIFS.
@@ 4.3 Use of Reading Order
Due to the unique feature of comics that the speaker is more likely to appear in the same or next frame that the text belongs to, we add frame information to the SGG model.
We train another Faster-RCNN model to detect the frame regions from the input image. To estimate the reading order of these frames, we follow the algorithm described in Kovanen and Ikuta’s study [17, 18]. As shown in Fig. 4, the outcome is a series of ordered frame bounding boxes $\mathbf{p}_{1},\mathbf{p}_{2},\dots$, where $\mathbf{p}_{i}\in\mathbb{R}^{4}$ is a bounding box of the $i$-th frame. We show the details of this process and results of frame detection and reading order estimation in the supplementary material.
@@
Here, $k_{i}$ represents the reading order of the frame $\mathbf{p}$ to which $\mathbf{b}_{i}$ belongs. The closer the reading orders of $\mathbf{b}_{i}$, $\mathbf{b}_{j}$ are, the higher the score of $w(i,j)$. Therefore, we can enhance prediction accuracy by incorporating frame orders and adjusting relationship scores accordingly.